Find the convective acceleration at the middle of a pipe which converges uniformly from 0.6 m diameter to 0.3 m diameter over 3 cm length. The rate of flow is 40 lit/s. If the rate of flow changes uniformly from 40 lit/s to 80 lit/s in 40 seconds, find the total acceleration at the middle of the pipe at 20th second.
1
Expert's answer
2020-12-15T12:01:53-0500

"D_1=0.6m, D_2=0.3m,L=3cm=0.03m,"

"Q=40L\/s=0.04m^3\/s,"

"Q_1=40l\/s=0.04m^3\/s,Q_2=80l\/s=0.08m^3\/s"

Case (i)

Flow is one dimensional and hence the velocity components "v=w=0"

Convective acceleration "=u(\partial u/\partial x)"

"A_1=\dfrac{\pi D_1^2}{4}=0.09\pi m^2\approx0.2827m^2"

"A_2=\dfrac{\pi D_2^2}{4}=0.0225\pi m^2\approx0.0707m^2"


"u_1=\dfrac{Q}{A_1}=\dfrac{0.04m^3/s}{0.09\pi m^2}\approx0.1415m/s"


"u_2=\dfrac{Q}{A_2}=\dfrac{0.04m^3/s}{0.0225\pi m^2}\approx0.5657m/s"

As the diameter changes uniformly, the velocity will also change uniformly. The velocity "u" at any distance "x" from inlet is given by


"u = u_1 +\dfrac{u_2 – u_1}{L} x"


"\dfrac{\partial u}{\partial x}=\dfrac{u_2-u_1}{L}"

"\text{Convective acceleration }=u\cdot \dfrac{u_2-u_1}{L}"

"x=0.015m"


"u = 0.1415m/s +\dfrac{0.5657m/s – 0.1415m/s}{0.03m} \cdot0.015m"

"=0.3536m\/s"

"\text{Convective acceleration }"




"=0.3536m/s\cdot \dfrac{0.5657m/s – 0.1415m/s}{0.03m}"

"=8.3355m\/s^2"

Case (ii)

Total acceleration = (convective + local ) acceleration at t =20 seconds

Rate of flow

"Q_{t=20}=Q_1+\dfrac{Q_2-Q_1}{40s}\cdot20s"

"=0.06m^3\/s"

"u_1=Q\/A_1=0.2122m\/s"

"u_2=Q\/A_2=0.8488m\/s"

The velocity "u" at any distance "x" from inlet is given by


"u = u_1 +\dfrac{u_2 – u_1}{L} x"

"u = 0.2122 +21.2207 x"

"\dfrac{\partial u}{\partial x}=21.2207s^{-1}""\text{Convective acceleration }=u\cdot \dfrac{\partial u}{\partial x}"

"=(0.2122 +21.2207 x)\cdot21.2207"


At "x=0.015m"


"(\text{Convective acceleration })_{x=0.015m}"

"=11.2578m\/s^2"


Local acceleration

Diameter at "x=0.015m" is given by 


"D = D_1 +\dfrac{D_2 – D_1}{L} x=0.45m"

"A=\dfrac{\pi D^2}{4}\approx0.1590m^2"


"u_1=Q_1\/A=0.2515m\/s"

"u_2=Q_2\/A=0.5030m\/s"

"\text{Rate of change of velocity}"

"=\dfrac{0.5030m/s-0.2515m/s}{40s}=0.0063m/s^2"



Total acceleration="=11.2578m\/s^2+0.0063m\/s^2=11.2641m\/s^2"